OpenStudy (anonymous):

The speed of light travels at a rate of 3.0 x 10^8 meters per second. If the Earth is 1.5 x 10^11 meters away from the sun, how long does it take for a ray of light to travel from the sun to the earth?

OpenStudy (anonymous):

\[\frac{1.5 x 10{}^{\wedge}11}{3.0 x 10{}^{\wedge}8}*\frac{1}{60}=8.33 \text{ minutes} \]

OpenStudy (paxpolaris):

\[D=r \times t\] \[\implies t= D \times \frac 1r= 1.5\cdot10^{11}m \times {\sec. \over 3.0\cdot10^8m}\]

OpenStudy (anonymous):

What is that in seconds?

OpenStudy (anonymous):

500 seconds

OpenStudy (paxpolaris):

\[\Large{\cancel{1.5}\cdot10^{11}\over \cancel{3.0}2\cdot10^8}= \frac12 \times 10^{11-8} = 1000/2=500\]

OpenStudy (anonymous):

Ok thank you